# Evolution With Purpose: Hierarchy-Informed Optimization of Whole-Brain Models

### v2

@@ Abstract.
copyright: noneconference: $c$s: Computing methodologies Genetic algorithmsccs: Mathematics of computing Optimization algorithmsccs: Computing methodologies Bio-inspired approachesccs: Computing methodologies Distributed algorithms
@@ 1. Introduction
Large-scale functional brain dynamics are commonly modeled as emergent phenomena arising from the interaction between anatomical connectivity and local neural physiology. Models such as neural mass, neural field, and Dynamic Mean Field (DMF) successfully reproduced empirical functional connectivity and dynamics at the level of resting-state networks (RSN)s (Honey et al., 2009; Deco et al., 2011; Breakspear, 2017; Maile et al., 2019). These models depend on many interacting hyperparameters that must be carefully tuned to match individual neuroimaging data. Evolutionary algorithms are well suited for this purpose because they explore complex, non-differentiable fitness landscapes without gradient information (Hansen, 2016; Miikkulainen, 2019; Pelikan, 2005). Prior work applying evolutionary strategies to DMF models demonstrated improved fits to resting-state data (Maile et al., 2019), but also highlighted that such fits often fail to generalize across subjects and may not predict behavior well. Similar issues of deception and overfitting arise in other real-world search domains. Although evolutionary computation (EC) techniques are well suited for large, high-dimensional, and deceptive search spaces—even at extreme scales involving thousands of states and billions of variables (Deb and Myburgh, 2017; Gomez and Miikkulainen, 1997; Miikkulainen, 2019; Shahrzad et al., 2018, 2019)—they can still converge to brittle or uninformative solutions in the absence of appropriate inductive biases.
@@
This work investigates a previously underexplored factor in evolutionary optimization of dynamical systems: the temporal order in which parameters are introduced during search. Drawing on cortical hierarchies, established in humans and primates as a principal gradient of cortical organization (Margulies et al., 2016), and incremental optimization insights (Gomez and Miikkulainen, 1997), heterogeneous optimization was compared with hierarchy-informed curriculum (HICO), reverse-phased curriculum, and shuffled curriculum. Although introducing heterogeneity substantially improved in-sample fitness, only flat (non-curricular) heterogeneous, HICO, and reverse-phased strategies significantly outperformed the homogeneous baseline ($p\approx 10^{-23}$$10^{-25}$). Marked differences emerged, however, under leave-one-out cross-subject evaluation using cohort-averaged parameters: flat heterogeneous and homogeneous baselines failed to generalize, whereas curriculum-based strategies remained robust. Among these, HICO yielded the most consistent and well-structured parameter solutions. Moreover, models based on HICO provided the strongest and most reliable predictions of cognitive and psychopathology targets spanning multiple domains, including fluid reasoning ability (as measured using the Penn Matrix Test (Moore et al., 2014)), behavioral problems related to internalization and externalization, as measured using the Achenbach Adult Self Report (Achenbach et al., 2017). These results indicate that incorporating domain-informed structure through evolutionary curricula fundamentally alters search dynamics, providing a principled mechanism for scaling high-dimensional brain models while yielding more stable, generalizable, and behaviorally informative solutions than flat optimization.
@@ 2.1. Whole-Brain Biophysical Models
A central goal of biophysical brain modeling is to explain how distributed functional dynamics emerge from the coupling of structural connectivity and local neural processes (Deco et al., 2011; Breakspear, 2017). Popular frameworks include neural mass, neural field, and Dynamic Mean Field (DMF) approaches. The resulting models can generate resting-state activity that reproduces key empirical signatures, including functional connectivity, dynamic functional connectivity, and metastability (Deco et al., 2013; Cabral et al., 2014).
A canonical study by Honey et al. (2009) showed that structural connectivity shapes—but does not fully determine—functional connectivity, motivating the use of computational models to explore this mapping. Subsequent work has characterized how factors such as excitation–inhibition balance, global coupling strength, conduction delays, and noise influence emergent functional patterns (Deco et al., 2011; Breakspear, 2017). The DMF model used in this paper is one such framework, balancing biological plausibility with computational tractability (Deco et al., 2014).
$2.$ Challenges in Parameter Optimization
@@ 2.3. Cortical Hierarchy and Training Curricula
The brain is typically parcellated into canonical RSNs) Yeo et al. (2011). More recently, Margulies et al. (2016) identified a dominant gradient ranging from unimodal (sensorimotor) to transmodal (default mode) RSNs, thereby providing a macroscale gradient of cortical organization.
@@ 2.6. Evolutionary Algorithms in Neuroscience
Evolutionary algorithms (EAs) maintain a population of candidate solutions that evolve under selection, mutation, and recombination (Holland, 1975). They are especially effective for high-dimensional, non-differentiable problems with complex fitness landscapes. Early applications demonstrated their value in fitting single-neuron conductance parameters (Vanier and Bower, 1999; Keren et al., 2005). Achard and De Schutter (2006) used evolutionary optimization to fit detailed Purkinje cell models, revealing high-dimensional manifolds of viable solutions. Druckmann et al. (2007) applied Multi-objective evolutionary algorithms to match multiple electrophysiological features simultaneously.
At larger scales, population-based evolutionary optimization has been used to tune RSNs and whole-brain models (Van Geit et al., 2016; Maile et al., 2019). While covariance-based strategies such as CMA-ES can be effective in moderate-dimensional settings (Hansen, 2016), they scale poorly as the number of parameters grows and impose strong assumptions on parameter correlations. In contrast, genetic-algorithm–style evolutionary computation offers greater scalability and architectural flexibility, making it well suited for high-dimensional whole-brain models and curriculum-guided parameter schedules. These insights motivate the present focus on introducing structured inductive biases into evolutionary optimization.
@@ 3.1. Data and Parcellation
Resting-state functional Magnetic Resonance Imaging (rs-fMRI) and diffusion-weighted imaging (DWI) data were obtained from 100 unrelated subjects from the Human Connectome Project (HCP). The cerebral cortex was parcellated into 400 regions using the Schaefer et al. (2017) atlas, with parcels assigned to seven canonical RSNs according to the Yeo et al. (2011) parcellation (Fig. 1). Subject-specific structural connectivity (SC) matrices were derived from tractography of the DWI data, and empirical functional connectivity (FC) matrices were computed as Pearson correlations of regional Blood-Oxygen-Level–Dependent (BOLD) rs-fMRI time series
@@ 3.2. Dynamic Mean Field Model
Each cortical region was modeled as a coupled excitatory–inhibitory neural mass using the Dynamic Mean Field (DMF) framework (Deco et al., 2011, 2013; Breakspear, 2017). The DMF model provides a mesoscale approximation to large populations of spiking neurons; it was shown to reproduce key properties of resting-state dynamics when driven by subject-specific structural connectivity (Deco et al., 2013; Murray et al., 2014).
Let $S_{i}(t)$ denote the excitatory synaptic gating variable for region ii. The excitatory population dynamics are governed by
@@
where $I_{b}$ is a background current, $w_{E}$ and $w_{I}$ denote local excitatory and inhibitory coupling strengths, $S^{I}_{i}(t)$ is the inhibitory gating variable, $G$ is the global coupling parameter, $\mathrm{SC}_{ij}$ is the structural connectivity weight from region $j$ to $i$, and $d_{ij}$ represents conduction delays. Canonical DMF constants were adopted from prior work (Deco et al., 2013), except for the parameters optimized in this study. See Appendix A.1 for the full specification of DMF variables and constants. The homogeneous model comprised 20 shared parameters applied uniformly across all regions. The heterogeneous formulation comprised 140 parameters, corresponding to seven RSNs with independent copies of the same 20 parameters. See Appendix A.2 for the full parameter specification. In both cases, the DMF equations define a stochastic dynamical system that generates region-wise neural activity time series whose statistical structure depends on the parameter vector $\Theta$.
$3.$ Simulation and Objective Function
@@
This FC-based objective follows standard practice in whole-brain modeling, where functional connectivity serves as a compact summary of long-range statistical dependencies induced by the underlying neural dynamics (Deco et al., 2013; Murray et al., 2014; Maile et al., 2019). Minimizing $\mathcal{L}(\Theta)$ therefore corresponds to calibrating the biophysical parameters of the DMF model so that its emergent RSN-level behavior reproduces the observed functional organization of the human brain.
@@ 3.4. Optimization Curricula
A single set of 20 parameters was used for all RSNs. Thus, this baseline assumes that biophysical dynamics is homogeneous throughout the brain.
@@
  • Phase I — Global Phase. This phase establishes the homogeneous baseline as the starting point. Further optimizaiton is thus restricted to a low-dimensional, dynamically coherent subspace, which reduces the unstable fixed points and improves Lyapunov stability (Deco et al., 2013).
@@ 3.5. Evolutionary Optimization Method
To ensure comparable search resolution across parameters with different physical units, each parameter was represented in a normalized coordinate in $[0,1]$ with 1000-step precision and mapped via an affine transform to its literature-supported valid range, as established in prior DMF modeling studies (Wong and Wang, 2006; Deco et al., 2013; Breakspear, 2017). This representation yields an effective discrete search space of $10^{60}$ possible configurations for the homogeneous model and $10^{420}$ for the heterogeneous formulation. This scale makes the evolutionary approach a good choice for optimizing these models.
@@
All experiments were parallelized using a hub-and-spoke distributed evolutionary architecture with four concurrent evolutionary engines (Shahrzad and Miikkulainen, 2023). In this setup, multiple worker spokes evaluate candidate solutions in parallel and synchronize with a central hub that maintains population state and applies selection and variation. This design increases throughput without changing the underlying optimization dynamics, and it ensures identical GA settings across homogeneous and heterogeneous curricula.
@@ 4.3. Geometry of Solutions
$4.$ Predicting Behavior
Behavioral relevance of optimized DMF parameters was assessed by predicting three behavioral targets from RSN-level parameter summaries: fluid reasoning ability (PMAT24_A_CR), behavioral problems related to internalization (inwardly directed affective and somatic symptoms (ASR_Intn_Raw)), and externalization (outwardly directed behavioral tendencies (ASR_Extn_Raw)). For each approach, parameters for the seven RSNs were averaged into a single 20 dimensional vector and ridge regression models were fit to each behavioral target separately. Predictive performance was then quantified using the coefficient of determination ($R^{2}$), which measures the proportion of behavioral variance explained by the model, relative to a baseline that predicts the target’s mean across subjects.
@@ 5. Discussion and Future Work
A natural next step is to scale this framework to larger and more diverse datasets. In particular, extending the analysis to the full Human Connectome Project (with $n\approx 1200$ subjects) will allow characterizing interindividual variability more precisely, thus making behavior predictions more accurate. In parallel, applying the proposed optimization framework to translational and clinical datasets will make it possible to study how interventions such as neuromodulation or pharmacological treatments alter the underlying dynamical regimes, providing a pathway for using whole-brain models to assess treatment-induced changes in large-scale brain organization.
@@ 6. Conclusion
This study demonstrated that the curricular structure of parameter optimization plays a critical role in generality and relevance of whole-brain biophysical models. Biologically motivated structure acts as a powerful inductive bias in the evolutionary optimization: it uncovers parameter regimes that generalize effectively across subjects, and can be used to predict behavior. Curriculum-informed evolutionary optimization thus provides a principled approach for scaling complex biophysical models. This framework offers a promising direction for future work in large-scale computational neuroscience and other high-dimensional scientific modeling domains.
@@ A.1. DMF State Variables and Fixed Constants
Each cortical region $i$ is modeled as a coupled excitatory–inhibitory neural population. The dynamical state is described by synaptic gating variables $S_{i}(t)$ (excitatory) and $S^{I}_{i}(t)$ (inhibitory), evolving according to the DMF equations defined in Section $2.$
The following constants were fixed across all experiments and adopted from prior DMF studies (Deco et al., 2013; Murray et al., 2014; Breakspear, 2017).
@@
The sigmoidal transfer function $H(\cdot)$ follows the standard DMF formulation (Deco et al., 2013), mapping total synaptic input to population firing rate.
@@ A.2. Optimized Parameters per RSN
The Dynamic Mean Field (DMF) model parameters optimized in this study follow the formulation of Deco et al. (2011, 2013); Murray et al. (2014). Each RSN is associated with an identical ordered block of 20 parameters.

### v4

@@ Abstract.
journalyear: 2026copyright: ccconference: Genetic and Evolutionary Computation Conference; July 13–17, 2026; San Jose, Costa Ricabooktitle: Genetic and Evolutionary Computation Conference (GECCO ’26), July 13–17, 2026, San Jose, Costa Ricadoi: 10.1145/3795095.3805080isbn: 979-8-4007-2487-9/2026/07ccs: Computing methodologies Genetic algorithmsccs: Mathematics of computing Optimization algorithmsccs: Computing methodologies Bio-inspired approachesccs: Computing methodologies Distributed algorithms
@@ 1. Introduction
Large-scale functional brain dynamics are commonly modeled as emergent phenomena arising from the interaction between anatomical connectivity and local neural physiology. Models such as neural mass, neural field, and Dynamic Mean Field (DMF) successfully reproduced empirical functional connectivity and dynamics at the level of resting-state networks (RSN)s (Honey et al., 2009; Deco et al., 2011; Breakspear, 2017; Maile et al., 2019). These models depend on many interacting hyperparameters that must be carefully tuned to match individual neuroimaging data. Evolutionary algorithms are well suited for this purpose because they explore complex, non-differentiable fitness landscapes without gradient information (Hansen, 2016; Miikkulainen, 2020; Pelikan, 2005; Demirtaş et al., 2019). Prior work applying evolutionary strategies to DMF models demonstrated improved fits to resting-state data (Maile et al., 2019), but also highlighted that such fits often fail to generalize across subjects and may not predict behavior well. Similar issues of deception and overfitting arise in other real-world search domains. Although evolutionary computation (EC) techniques are well suited for large, high-dimensional, and deceptive search spaces—even at extreme scales involving thousands of states and billions of variables (Deb and Myburgh, 2017; Gomez and Miikkulainen, 1997; Miikkulainen, 2020; Shahrzad et al., 2018, 2019)—they can still converge to brittle or uninformative solutions in the absence of appropriate inductive biases.
@@
This work investigates a previously underexplored factor in evolutionary optimization of dynamical systems: the temporal order in which parameters are introduced during search. Drawing on cortical hierarchies, established in humans and primates as a principal gradient of cortical organization (Margulies et al., 2016), and incremental optimization insights (Gomez and Miikkulainen, 1997), heterogeneous optimization was compared with hierarchy-informed curriculum (HICO), reverse-phased curriculum, and shuffled curriculum. Although introducing heterogeneity substantially improved in-sample fitness, only flat (non-curricular) heterogeneous, HICO, and reverse-phased strategies significantly outperformed the homogeneous baseline ($p\approx 10^{-23}$$10^{-25}$). Marked differences emerged, however, under leave-one-out cross-subject evaluation using cohort-averaged parameters: flat heterogeneous and homogeneous baselines failed to generalize, whereas curriculum-based strategies remained robust. Among these, HICO yielded the most consistent and well-structured parameter solutions. Moreover, models based on HICO provided the strongest and most reliable predictions of cognitive and psychopathology targets spanning multiple domains, including fluid reasoning ability (as measured using the Penn Matrix Test (Moore et al., 2015)), behavioral problems related to internalization and externalization, as measured using the Achenbach Adult Self Report (Achenbach et al., 2017). These results indicate that incorporating domain-informed structure through evolutionary curricula fundamentally alters search dynamics, providing a principled mechanism for scaling high-dimensional brain models while yielding more stable, generalizable, and behaviorally informative solutions than flat optimization.
@@ 2.1. Whole-Brain Biophysical Models
A central goal of biophysical brain modeling is to explain how distributed functional dynamics emerge from the coupling of structural connectivity and local neural processes (Deco et al., 2011; Breakspear, 2017). Popular frameworks include neural mass, neural field, and Dynamic Mean Field (DMF) approaches. The resulting models can generate resting-state activity that reproduces key empirical signatures, including functional connectivity, dynamic functional connectivity, and metastability (Deco et al., 2009; Cabral et al., 2014).
A canonical study by Honey et al. (2009) showed that structural connectivity shapes—but does not fully determine—functional connectivity, motivating the use of computational models to explore this mapping. Subsequent work has characterized how factors such as excitation–inhibition balance, global coupling strength, conduction delays, and noise influence emergent functional patterns (Deco et al., 2011; Breakspear, 2017). The DMF model used in this paper is one such framework, balancing biological plausibility with computational tractability (Deco et al., 2013).
2.2. Challenges in Parameter Optimization
@@ 2.3. Cortical Hierarchy and Training Curricula
For analysis, the brain is often parcellated into RSNs, i.e. functional subdivisions that are well-established in the literature (Yeo et al., 2011). Recently, Margulies et al. (2016) identified a dominant gradient ranging from unimodal (sensorimotor) to transmodal (default mode) RSNs, thereby providing a macroscale order for cortical organization.
@@ 2.6. Evolutionary Algorithms in Neuroscience
Evolutionary algorithms (EAs) maintain a population of candidate solutions that evolve under selection, mutation, and recombination (Holland, 1992). They are especially effective for high-dimensional, non-differentiable problems with complex fitness landscapes. Early applications demonstrated their value in fitting single-neuron conductance parameters (Vanier and Bower, 1999; Keren et al., 2005). Subsequently they were used to fit detailed Purkinje cell models, revealing high-dimensional manifolds of viable solutions (Achard and De Schutter, 2006). Further, using the multi-objective evolutionary approach, it was possible to match multiple electrophysiological features simultaneously (Druckmann et al., 2007). At larger scales, population-based evolutionary optimization has been used to tune RSNs and whole-brain models (Van Geit et al., 2016; Maile et al., 2019).
Among the many variants of EAs, covariance-based strategies such as CMA-ES have proven highly effective in moderate dimensional settings (Hansen, 2016). However, they are likely to be less effective in the high-dimensional spaces such as those resulting from neuroscience applications. For this reason, this paper relies on the traditional genetic algorithms: they are scalable even with future extensions to large-scale models. Further, they are flexible, and can thus readily accommodate structured interventions such as the curriculum-based parameter scheduling. In any case, the experiments in this paper do not rely on any optimization method in particular; the goal is to evaluate the effect of curriculum as an inductive bias in general.
@@ 3.1. Data and Parcellation
Resting-state functional Magnetic Resonance Imaging (rs-fMRI) and diffusion-weighted imaging (DWI) data were obtained from 100 unrelated subjects from the Human Connectome Project (HCP). The cerebral cortex was parcellated into 400 regions using the Schaefer et al. (2018) atlas, with parcels assigned to seven canonical RSNs according to the Yeo et al. (2011) parcellation (Fig. 1). Subject-specific structural connectivity (SC) matrices were derived from tractography of the DWI data, and empirical functional connectivity (FC) matrices were computed as Pearson correlations of regional Blood-Oxygen-Level–Dependent (BOLD) rs-fMRI time series.
@@ 3.2. Dynamic Mean Field Model
Each cortical region was modeled as a coupled excitatory–inhibitory neural mass using the Dynamic Mean Field (DMF) framework (Deco et al., 2011, 2009; Breakspear, 2017). The DMF model provides a mesoscale approximation to large populations of spiking neurons; it was shown to reproduce key properties of resting-state dynamics when driven by subject-specific structural connectivity (Deco et al., 2009; Murray et al., 2014).
Let $S_{i}(t)$ denote the excitatory synaptic gating variable for region $i$. The excitatory population dynamics are governed by
@@
where $I_{b}$ is a background current, $w_{E}$ and $w_{I}$ denote local excitatory and inhibitory coupling strengths, $S^{I}_{i}(t)$ is the inhibitory gating variable, $G$ is the global coupling parameter, $\mathrm{SC}_{ij}$ is the structural connectivity weight from region $j$ to $i$, and $d_{ij}$ represents conduction delays. Canonical DMF constants were adopted from prior work (Deco et al., 2009), except for the parameters optimized in this study. See Appendix A.1 for the full specification of DMF variables and constants. The homogeneous model comprised 20 shared parameters applied uniformly across all regions. The heterogeneous formulation comprised 140 parameters, corresponding to seven RSNs with independent copies of the same 20 parameters. See Appendix A.2 for the full parameter specification. In both cases, the DMF equations define a stochastic dynamical system that generates region-wise neural activity time series whose statistical structure depends on the parameter vector $\Theta$.
3.3. Simulation and Objective Function
@@
This FC-based objective follows standard practice in whole-brain modeling, where functional connectivity serves as a compact summary of long-range statistical dependencies induced by the underlying neural dynamics (Deco et al., 2009; Murray et al., 2014; Maile et al., 2019). Minimizing $\mathcal{L}(\Theta)$ therefore corresponds to calibrating the biophysical parameters of the DMF model so that its emergent RSN-level behavior reproduces the observed functional organization of the human brain.
@@ 3.4. Optimization Curricula
A single set of 20 parameters was used for all RSNs. Thus, this baseline assumes that biophysical dynamics are homogeneous throughout the brain.
@@
  • Phase I — Global Phase. This phase establishes the homogeneous baseline as the starting point. Further optimizaiton is thus restricted to a low-dimensional, dynamically coherent subspace, which reduces the unstable fixed points and improves Lyapunov stability (Deco et al., 2009).
@@ 3.5. Evolutionary Optimization Method
To ensure comparable search resolution across parameters with different physical units, each parameter was represented in a normalized coordinate in $[0,1]$ with 1000-step precision and mapped via an affine transform to its literature-supported valid range, as established in prior DMF modeling studies (Wong and Wang, 2006; Deco et al., 2009; Breakspear, 2017). This representation yields an effective discrete search space of $10^{60}$ possible configurations for the homogeneous model and $10^{420}$ for the heterogeneous formulation. This scale makes the evolutionary approach a good choice for optimizing these models.
@@
All experiments were parallelized using a hub-and-spoke distributed evolutionary architecture with four concurrent evolutionary engines (Shahrzad and Miikkulainen, 2023). In this setup, multiple worker spokes evaluate candidate solutions in parallel and synchronize with a central hub that maintains population state and applies selection and variation. This design increases throughput without changing the underlying optimization dynamics, and ensures identical GA settings across all methods. It thus enables a controlled comparison that isolates the impact of curriculum on the optimization process.
@@ 4.3. Geometry of Solutions
4.4. Predicting Behavior
Behavioral relevance of optimized DMF parameters was assessed by predicting three behavioral targets from RSN-level parameter summaries: fluid reasoning ability (PMAT24_A_CR) and behavioral problems related to either internalization (inwardly directed affective and somatic symptoms, ASR_Intn_Raw) or externalization (outwardly directed behavioral tendencies, ASR_Extn_Raw). For each approach, parameters for the seven RSNs were averaged into a single 20 dimensional vector and ridge regression models were fit to each behavioral target separately. Predictive performance was then quantified using the coefficient of determination ($R^{2}$), which measures the proportion of behavioral variance explained by the model, relative to a baseline that predicts the target’s mean across subjects.
@@ 5. Discussion and Future Work
A natural next step is to scale this framework to larger and more diverse datasets. In particular, extending the analysis to the full Human Connectome Project (with $n\approx 1200$ subjects) will allow characterizing inter-individual variability more precisely, thus making behavior predictions more accurate. In parallel, applying the proposed optimization framework to translational and clinical datasets will make it possible to study how interventions such as neuromodulation or pharmacological treatments alter the underlying dynamical regimes, providing a pathway for using whole-brain models to assess treatment-induced changes in large-scale brain organization.
@@ 6. Conclusion
This study demonstrated that the curricular structure of parameter optimization plays a crucial role in generality and relevance of whole-brain biophysical models. Biologically motivated structure acts as a powerful inductive bias in the evolutionary optimization: it uncovers parameter regimes that generalize effectively across subjects, and can be used to predict behavior. Curriculum-informed evolutionary optimization thus provides a principled approach for scaling complex biophysical models. This framework offers a promising direction for future work in large-scale computational neuroscience and other high-dimensional scientific modeling domains.
@@ A.1. DMF State Variables and Fixed Constants
Each cortical region $i$ is modeled as a coupled excitatory–inhibitory neural population. The dynamical state is described by synaptic gating variables $S_{i}(t)$ (excitatory) and $S^{I}_{i}(t)$ (inhibitory), evolving according to the DMF equations defined in Section 2.2.
The following constants were fixed across all experiments and adopted from prior DMF studies (Deco et al., 2009; Murray et al., 2014; Breakspear, 2017).
@@
The sigmoidal transfer function $H(\cdot)$ follows the standard DMF formulation (Deco et al., 2009), mapping total synaptic input to population firing rate.
@@ A.2. Optimized Parameters per RSN
The Dynamic Mean Field (DMF) model parameters optimized in this study follow the formulation of Deco et al. (2011, 2009); Murray et al. (2014). Each RSN is associated with an identical ordered block of 20 parameters.